# Holomorphic bootstrap for RCFT: signs and bounds for quasi-characters

### v1

@@ 5.4 Non-identity character
Now if we can prove $\alpha(R)\gtrsim R$ then we are done. This in turn can be done if there is a real solution to $\alpha(R)=R$ with $\alpha(R)$ given by Eq. (5.53). Numerically we find there are three real solutions, $\simeq 0.33,41.64,52.59$, with $\alpha(R)>R$ between the latter two. Thus for $R$ anywhere in the range $5242<R<$ 52, the desired inequality holds.
Thus we have proved that $b_{n}>Rb_{n-1}$ for all $n<2|M|$ which implies that $b_{n}>0$ and are monotonic. This in turn shows that $a_{n}$ are alternating up to $n=2|M|-1$.
Next, we see why this result fails above $n=2|M|-1$. This is because $f_{M}(n,i)>0$ for $n=2|M|$ and hence in Eq. (5.47), the leading term becomes negative and hence $b_{2M}<0$ implying $a_{2M}<0$.
$2|M|\leq n<15|M|^{2}$
@@
Note that, for $i\equiv n-2|M|+2k+1$ with $10\leq k\leq|M|-$ 1, we have:
@@
where we have used that $a_{k}>0$ for all $20\leq k\leq n-$ 2. Hence, proved.
@@ Acknowledgements
The work of A.D. is supported by the STFC Consolidated Grant ST/T000600/1 – “Particle Theory at the Higgs Centre”. He gratefully acknowledges King’s College London and the organisers of ICFT 2025: UK Meetings on Integrable and Conformal Field Theory – especially Gerard Watts, where preliminary results of this work were presented. He also thanks Jishu Das, Faisal Karimi, Anatoly Konechny, Alexander Radcliffe, Naveen Balaji Umasankar and Gerard Watts for their insightful discussions on RCFTs and modular forms. S.M. is supported by the Raja Ramanna Chair of the Department of Atomic Energy, Government of India. He would also like to thank Yuji Tachikawa and the Institute for the Mathematics and Physics of the Universe (IPMU), University of Tokyo, for their warm and generous hospitality during the period that this work was being completed.
@@
Notice that there are no quasi-characters in this range $-4<c<0$. It then follows from Eq. (A.3) that precisely one of the two exponents must be positive and the other negative. We now re-organise the above quasi-characters in a slightly different way from [1] that makes it easier to understand the construction of admissible characters. To do this, let us define the following four series at $\ell=0$:
@@
We have restricted the range of $n$ to cover exactly half the cases in Eq. (A.1). However we now take $n\in\mathbb{Z}$, so it runs over all positive and negative integers. Then the remaining half of the solution set, that we excluded by so restricting $p$, arises at negative values of $p$ but with the role of the identity and non-identity solution reversed. For example, in the $A_{1}$ series we have excluded $p=2$ which would have $c=13$, however we now allow $p=-3$ which corresponds to $c=-17$. This is exactly $-c-4$ for $c=13$, which means the $c=13$ solution has re-appeared with an exchange of the two components. Notice that when $c$ is negative, $h$ is also negative and therefore the exponent of the identity character is less singular than that of the non-identity character.

### v3

@@ 5.4 Non-identity character
Now if we can prove $\alpha(R)\gtrsim R$ then we are done. This in turn can be done if there is a real solution to $\alpha(R)=R$ with $\alpha(R)$ given by Eq. (5.53). Numerically we find there are three real solutions, $\simeq 0.33,41.64,52.59$, with $\alpha(R)>R$ between the latter two. Thus for $R$ anywhere in the range $42<R<52$, the desired inequality holds.
Thus we have proved that $b_{n}>Rb_{n-1}$ for all $n<2|M|$ which implies that $b_{n}>0$ and are monotonic. This in turn shows that $a_{n}$ are alternating up to n=2|M|1n=2|M|-1.
Next, we see why this result fails above n=2|M|1n=2|M|-1. This is because $f_{M}(n,i)>0$ for $n=2|M|$ and hence in Eq. (5.47), the leading term becomes negative and hence $b_{2M}<0$ implying $a_{2M}<0$.
$\bm{2|M|}\leq\bm{n}<\bm{15|M|^{2}}$
@@
Note that, for $i\equiv n-2|M|+2k+1$ with $0\leq k\leq|M|-1$, we have:
@@
where we have used that $a_{k}>0$ for all $0\leq k\leq n-2$. Hence, proved.
@@ Acknowledgements
The work of A.D. is supported by the STFC Consolidated Grant ST/T000600/1 – “Particle Theory at the Higgs Centre”. He gratefully acknowledges King’s College London and the organisers of ICFT 2025: UK Meetings on Integrable and Conformal Field Theory – especially Gerard Watts, where preliminary results of this work were presented. He also thanks Jishu Das, Faisal Karimi, Anatoly Konechny, Alexander Radcliffe, Jagannath Santara, Naveen Balaji Umasankar and Gerard Watts for their insightful discussions on RCFTs and modular forms. S.M. is supported by the Raja Ramanna Chair of the Department of Atomic Energy, Government of India. He would also like to thank Yuji Tachikawa and the Institute for the Mathematics and Physics of the Universe (IPMU), University of Tokyo, for their warm and generous hospitality during the period that this work was being completed.
@@
Notice that there are no quasi-characters in this range 4<c<0-4<c<0. It then follows from Eq. (A.3) that precisely one of the two exponents must be positive and the other negative. We now re-organise the above quasi-characters in a slightly different way from [1] that makes it easier to understand the construction of admissible characters. To do this, let us define the following four series at $\ell=0$:
@@
We have restricted the range of $n$ to cover exactly half the cases in Eq. (A.1). However we now take $n\in\mathbb{Z}$, so it runs over all positive and negative integers. Then the remaining half of the solution set, that we excluded by so restricting $p$, arises at negative values of $p$ but with the role of the identity and non-identity solution reversed. For example, in the $A_{1}$ series we have excluded $p=2$ which would have $c=13$, however we now allow p=3p=-3 which corresponds to c=17c=-17. This is exactly c4-c-4 for $c=13$, which means the $c=13$ solution has re-appeared with an exchange of the two components. Notice that when $c$ is negative, $h$ is also negative and therefore the exponent of the identity character is less singular than that of the non-identity character.
@@
Appendix B Exponential expression for exact $a_{n}$
In this appendix, we demonstrate that for the range $1\leq n\leq 2M$, the Fourier coefficients $a_{n}$, whose exact expression is given in equation Eq. (4.7), can be written as a product of $P_{M}(n)$ (see Eq. (4.12)) and an exponential function. The exponential will be a function of $n$, $M$, and a real parameter $\beta>0$. For simplicity, we present the argument for the identity character with M>0M>0191919The reasoning can be extended to the non-identity character and to cases with $M<0$ as well..
Let us take $n\sim{\cal O}(M)$ and $n\leq 2M$. We start by writing $a_{n}^{(2)}$ which is the part of $a_{n}$ (as in Eq. (4.7)) with terms involving only $g(k,2)$ factors and its products,
$a_{n}^{(2)}$ $\equiv P_{M}(n)\left[1+\sum_{k_{1}=1}^{n-1}g(k_{1},2)+\sum_{k_{1}=1}^{n-3}\sum_{k_{2}=k_{1}+2}^{n-1}g(k_{1},2)g(k_{2},2)\right$.
$+\sum_{k_{1}=1}^{n-5}\sum_{k_{2}=k_{1}+2}^{n-3}\sum_{k_{3}=k_{2}+2}^{n-1}g(k_{1},2)g(k_{2},2)g(k_{3},2)+\ldots$
$\left.+\sum_{k_{1}=1}^{n-(2i-1)}\sum_{k_{2}=k_{1}+2}^{n-(2i-3)}\ldots\sum_{k_{i}=k_{i-1}+2}^{n-1}g(k_{1},2)g(k_{2},2)\ldots g(k_{i},2)+\ldots\right$.
$\left.+\sum_{k_{1}=1}^{n-(2i_{\text{max}}-1)}\sum_{k_{2}=k_{1}+2}^{n-(2i_{\text{max}}-3)}\ldots\sum_{k_{i_{\text{max}}}=k_{i_{\text{max}}-1}+2}^{n-1}g(k_{1},2)g(k_{2},2)\ldots g(k_{i_{\text{max}}},2)\right]$ (B.1)
with $i_{\text{max}}=\left\lfloor\frac{n}{2}\right\rfloor$. Thus, the above expression has $i_{\text{max}}+1$ terms.
Now note that $a_{n}^{(2)}$ above is the leading part of $a_{n}$ at every order since it only contains $g(k,p)$ with $p=2$. Here we are omitting all $p\geq 3$ factors. Now we can estimate $a_{n}^{(2)}$ with:
$a_{n}^{\text{exp}}=P_{M}(n)\exp\left(\sum_{k_{1}=1}^{n-1}g(k_{1},2)\right)\equiv P_{M}(n)\exp\left(s^{(2)}\right).$ (B.2)
With this now let us compute quadratic error. At quadratic order, we define the error as:
$e^{(2)}\equiv\frac{(s^{(2)})^{2}}{2}-\sum_{k_{1}=1}^{n-3}\sum_{k_{2}=k_{1}+2}^{n-1}g(k_{1},2)g(k_{2},2),$ (B.3)
which after a bit of algebra, becomes:
$e^{(2)}=\frac{1}{2}\sum_{k=1}^{n-1}g(k,2)^{2}+\sum_{k=1}^{n-2}g(k,2)g(k+1,2),$ (B.4)
which shows that $e^{(2)}>0$.
From the inductive proof, we know that $P_{M}(n)$ as defined in Eq. (4.12) and $a_{n}$ have the same sign. Now let us define below,
$\tilde{r}(\beta,n,M)\equiv\frac{\left(\frac{a_{n}}{P_{M}(n)}\right)-\exp\left(s^{(2)}-\beta e^{(2)}\right)}{\left(\frac{a_{n}}{P_{M}(n)}\right)}.$ (B.5)
where $\beta\geq 0$ is an arbitrary real parameter for now  [^20] One way to understand $\beta$ is the following. Note that we are only considering the quadratic error at this point that is $e^{(2)}$. However, there will also be cubic and higher-order errors coming from exponentiating $s^{(2)}$ alone and comparing that expression to $a_{n}^{(2)}$. The true expression will be a result of exponentiating terms involving $g(k,3)$, $g(k,4)$, $\ldots$, etc., – that is all the single sum terms and not just $g(k,2)$. All of these higher-order error effects can be absorbed into a renormalised parameter and that is precisely the role of $\beta$ in Eq. (B.5)..
Now let us take $n=\alpha M$ with some fixed $\alpha$ in the range: $1<\alpha\leq 2$ in Eq. (B.5). Let us argue that $\tilde{r}(\beta,M)<0$ for $\beta=0$. We begin by claiming: $|a_{n}^{\text{exp}}|>|a_{n}^{(2)}|$. One way to see this is that when we expand the exponential in $a_{n}^{\text{exp}}$ then at each order the nested sums in $|a_{n}^{\text{exp}}|$ go from $1$ to n1n-1 but the upper limit of the nested sums in $a_{n}^{(2)}$ keep on decreasing by $2$ (as can be seen in Eq. (B.3)). This is exactly why $e^{(2)}>0$.
Furthermore, $a_{n}^{(2)}$ has a finite set of terms, that is, $i_{\text{max}}+1$ terms, to be precise. However, $a_{n}^{\text{exp}}$ has an infinite set of terms upon expanding the exponential. Thus, $|a_{n}^{\text{exp}}|>|a_{n}^{(2)}|$.
One might think including $a_{n}^{(3)}$ – which contains terms $g(k,3)$ and its products; $a_{n}^{(4)},\ldots$, etc., may lead to $|a_{n}|>|a_{n}^{\text{exp}}|$ but this won’t happen since we already know that these are sub-leading w.r.t. $a_{n}^{(2)}$ (see Eq. (4.37)). Hence, we have argued that: $|a_{n}^{\text{exp}}|>|a_{n}|$.
From the above discussion, we see that $\tilde{r}(\beta,M)<0$ for $\beta=0$. Now, note that, if we take a large enough $\beta>0$ then we can make $\tilde{r}(\beta,M)>0$. Since $\tilde{r}(\beta,M)$ for a fixed $M$ and as a function of $\beta$ is continuous in $\beta$, by the Intermediate Value Theorem, there will exist a $\beta\in(0,\infty)$, say $\beta^{*}>0$, such that $\tilde{r}(\beta^{*},M)=0$. Then, we will have  [^21] Numerically, we find that for $j=1$ and n=2M1n=2M-1 with large $2\leq M\leq 200$, $\beta^{*}\in(0,1)$. We conclude this by noting that for $\beta=0$, $\tilde{r}(0,M)<0$ and $|\tilde{r}(0,M)|$ increases with $M$. Additionally, for $\beta=1$, $\tilde{r}(1,M)>0$ and $|\tilde{r}(1,M)|$ again increases with $M$.:
$a_{n}=P_{M}(n)\exp\left(s^{(2)}-\beta^{*}e^{(2)}\right)$ (B.6)
with $n=\alpha M$ and for some fixed $\alpha$ in the range: $1<\alpha\leq 2$. Note that, here $a_{n}$ is exact and not approximate. Thus, Eq. (B.6) is the exponential expression in terms of $M$ and $\beta^{*}$ we were after. This shows that $a_{n}$ as given in Eq. (4.7) nicely exponentiates and further adds to the fact that the sign of $a_{n}$ is same as the sign of $P_{M}(n)$ for $1\leq n\leq 2M$.